OpenStudy (anonymous):

Can you help me solve this question average movie prices in the United States are in general lower than in other countries. It would cost $79.50 to buy three tickets in Japan plus two tickets in Switzerland. Three tickets in Switzerland plus two tickets in Japan would cost $76.20. How much does an average movies ticket cost in each of these countries?

OpenStudy (tkhunny):

Half the battle is in the definition. Write clear and concise defintiions. S = Movie price in Switzerland J = Movie Price in Japan. There, now you can translate the given information. "It would cost $79.50 to buy three tickets in Japan plus two tickets in Switzerland" "Three tickets in Switzerland plus two tickets in Japan would cost $76.20"

OpenStudy (anonymous):

I got this part tkhunny but do I divide and multiply later.

OpenStudy (tkhunny):

You got what part? There has yet to be any translation! Words "It would cost $79.50 to buy three tickets in Japan plus two tickets in Switzerland" Translation Using the Given Definitions $79.50 = 3*J + 2*S Stare at that until it makes sense. You do the other one. Words "Three tickets in Switzerland plus two tickets in Japan would cost $76.20" Translation Using the Given Definitions ??

OpenStudy (anonymous):

Okay tkhunny i got $76.20=3*76.20+2*79.50. Now what would be the next step. Please help.

OpenStudy (anonymous):

Thats okay I failed the test anyway thanks for the help.

OpenStudy (tkhunny):

This is why we write definitions. $76.20 and $79.50 are odd animals that are comprised of various numbers of various tickets. This is why we write definitions, so we will not forget what we mean. Words "It would cost $79.50 to buy three tickets in Japan plus two tickets in Switzerland" Translation Using the Given Definitions $79.50 = 3*J + 2*S Stare at that until it makes sense. Words "Three tickets in Switzerland plus two tickets in Japan would cost $76.20" Translation Using the Given Definitions $76.20 = 3*S + 2*J The challenge, now, is to find S and J. You have two equations to help do that, so it looks promising.